# Beyond Silent Letters: Amplifying LLMs in Emotion Recognition with Vocal Nuances

### v3

@@ Abstract
Emotion recognition in speech is a challenging multimodal task that requires understanding both verbal content and vocal nuances. This paper introduces a novel approach to emotion detection using Large Language Models (LLMs), which have demonstrated exceptional capabilities in natural language understanding. To overcome the inherent limitation of LLMs in processing audio inputs, we propose SpeechCueLLM, a method that translates speech characteristics into natural language descriptions, allowing LLMs to perform multimodal emotion analysis via text prompts without any architectural changes. Our method is minimal yet impactful, outperforming baseline models that require structural modifications. We evaluate SpeechCueLLM on two datasets: IEMOCAP and MELD, showing significant improvements in emotion recognition accuracy, particularly for high-quality audio data. We also explore the effectiveness of various feature representations and fine-tuning strategies for different LLMs. Our experiments demonstrate that incorporating speech descriptions yields a more than 2% increase in the average weighted F1 score on IEMOCAP (from 70.111% to 72.596%). The source code will be made available on GitHub.
@@ 2.1 Multimodal Emotion Models
Exsiting multimodal methods have been proposed to combine audio and text modalities for emotion recognition Zadeh et al. (2017); Wu et al. (2024); Meng et al. (2024); Tsai et al. (2019); Goncalves and Busso (2022); Ma et al. (2024a). Many of these works utilize transformer encoder structures to fuse multimodal signals through cross-modality attention mechanisms Tsai et al. (2019). Additionally, some models employ variations of Graph Neural Networks (GNNs) to capture multimodal semantic features and enhance emotion recognition accuracy Meng et al. (2024). These methods focus on model structure that captures and learns from multimodal signals holistically.
@@ 2.3 Speech Incorporation in LLMs
Recent researchhas also explored the application of LLMs in Automatic Speech Recognition (ASR)Bai et al. (2024); Baskar et al. (2024); Ma et al. (2024b) , demonstrating the versatility of LLMs in speech-processing tasks. Further extending the capabilities of LLMs in speech processing, AudioChatLlama Fathullah et al. (2024) and Audio-LLM Li et al. (2024) presents end-to-end model that integrates general-purpose speech abilities into the Llama-2 model.
@@ 4.4.3 How Do LLMs Differ in Performance?
Our results show consistent performance improvements across all models when either speech descriptions or impressions are integrated, underscoring the value of audio-derived features in emotion recognition tasks. In most cases, speech descriptions slightly outperform speech impressions, though the differences are often marginal. This suggests that more objective, feature-based descriptions may be more reliable than hard-coded speech impressions for emotion detection tasks.
@@
The substantial performance gap between the zero-shot Claude model and the fine-tuned models highlights the importance of task-specific fine-tuning for emotion recognition. However, even in a zero-shot setting, Claude benefits from the inclusion of speech features, further emphasizing the value of multimodal inputs in emotion detection.
@@ A.2 Evaluation of Speech Features Using Simple Machine Learning Models
To evaluate the effectiveness of our extracted speech features in capturing emotion-related information, we conducted experiments using various machine learning models on both numerical and categorical representations of these features. Tables 6 and 7 present the performance of Random Forest (RF), Support Vector Machine (SVM), and Multi-Layer Perceptron (MLP) models on the IEMOCAP and MELD datasets.

### v4

@@ Abstract
Emotion recognition in speech is a challenging multimodal task that requires understanding both verbal content and vocal nuances. This paper introduces a novel approach to emotion detection using Large Language Models (LLMs), which have demonstrated exceptional capabilities in natural language understanding. To overcome the inherent limitation of LLMs in processing audio inputs, we propose SpeechCueLLM, a method that translates speech characteristics into natural language descriptions, allowing LLMs to perform multimodal emotion analysis via text prompts without any architectural changes. Our method is minimal yet impactful, outperforming baseline models that require structural modifications. We evaluate SpeechCueLLM on two datasets: IEMOCAP and MELD, showing significant improvements in emotion recognition accuracy, particularly for high-quality audio data. We also explore the effectiveness of various feature representations and fine-tuning strategies for different LLMs. Our experiments reveal that incorporating speech descriptions leads to an improvement of nearly 10 points in the F1 score under the zero-shot setting and over 2.5 points under the LoRA setting on the IEMOCAP dataset. The source code will be made available on GitHub.
@@ 2.1 Multimodal Emotion Models
Exsiting multimodal methods have been proposed to combine audio and text modalities for emotion recognition Zadeh et al. (2017); Wu et al. (2024); Meng et al. (2024a); Tsai et al. (2019); Goncalves and Busso (2022); 10109845. Many of these works utilize transformer encoder structures to fuse multimodal signals through cross-modality attention mechanisms Tsai et al. (2019). Additionally, some models employ variations of Graph Neural Networks (GNNs) to capture multimodal semantic features and enhance emotion recognition accuracy Meng et al. (2024a). These methods focus on model structure that captures and learns from multimodal signals holistically.
@@ 2.3 Speech Incorporation in LLMs
Recent researchhas also explored the application of LLMs in Automatic Speech Recognition (ASR)Bai et al. (2024); Baskar et al. (2024); Ma et al. (2024b) , demonstrating the versatility of LLMs in speech-processing tasks. Further extending the capabilities of LLMs in speech processing, AudioChatLlama Fathullah et al. (2024) and Audio-LLM Li et al. (2024a) presents end-to-end model that integrates general-purpose speech abilities into the Llama-2 model.
@@ 4.3.1 Comparison with the Baselines
4.3.2 Comparison with the SOTAs
Method Modality IEMOCAP F1 Score MELD F1 Score
SDT Ma et al. (2024a) T, A, V 74.08 66.6
CFN-ESA Li et al. (2024b) T, A, V 71.04 66.70
GS-MCC Meng et al. (2024b) T, A, V 73.9 69.0
ELR-GNN Shou et al. (2024a) T, A, V 70.9 69.9
MPT-HCL Zou et al. (2023) T, A, V 72.51 65.02
TELME Yun et al. (2024) T, A, V 70.48 67.37
Mamba-like Model Shou et al. (2024b) T, A, V 73.3 67.6
SpeechCueLLM (ours) T, A 72.6 67.6
Table 2: Performance comparison of methods on the IEMOCAP and MELD datasets. Modality: T (Text), A (Audio), V (Video).
We collected results from the best-performing multimodal models in the recent two years for comparison. As shown in Table 2 Despite using only two modalities (Text and Audio), SpeechCueLLM’s performance is on par with SOTA models that utilize three modalities (Text, Audio, and Video) for both datasets. Our LLM-based approach also provides more interpretability for both modalities. Moreover, SpeechCueLLM requires only 12.8 million trainable parameters, which is less than two linear projection layers in the projection-based baseline, highlighting its efficiency.
@@ 4.4.3 How Do LLMs Differ in Performance?
The substantial performance gap between the zero-shot model and the fine-tuned models highlights the importance of task-specific fine-tuning for emotion recognition. However, even in a zero-shot setting, the models benefit from the inclusion of speech features, further emphasizing the value of multimodal inputs in emotion detection.
We observed significant improvements after applying speech descriptions in the zero-shot setting, with almost a 10-point increase in F1 scores for LLaMA-3-8b. This signifies the effectiveness of SpeechCueLLM and expands the use cases for applications that aim for fast deployment or do not have the resources for fine-tuning. The speech impressions also consistently improve upon the speech descriptions in the zero-shot setting for both models.
In the LoRA setting, our results show consistent performance improvements across all models when either speech descriptions or impressions are integrated, underscoring the value of audio-derived features in emotion recognition tasks. In most cases, speech descriptions slightly outperform speech impressions, though the differences are often marginal. This suggests that more objective, feature-based descriptions may be more reliable than hard-coded speech impressions for emotion detection tasks.
@@
4.4.4 Why SpeechCueLLM Gets Better Scores?
As highlighted in Table 9, our analysis across three independent runs demonstrates consistent improvements in the F1-score for all emotion classes when leveraging SpeechCueLLM. This underscores the robustness and reliability of the SpeechCueLLM approach in enhancing emotion recognition across diverse emotional categories.
The confusion matrices in Table 10 and Table 11 reveal a notable improvement in the accurate classification of ’Neutral’ emotions, increasing from 74.80% without speech features to 79.81% with SpeechCueLLM. While other emotion classes exhibit incremental gains, this substantial improvement in ’Neutral’ classification highlights SpeechCueLLM’s ability to discern the subtle nuances of emotionally subdued or calm utterances. This suggests that the model effectively captures the distinct characteristics of neutral expressions, particularly in scenarios where speakers convey emotions with minimal intensity or through a calm demeanor.
@@ A.2 Evaluation of Speech Features Using Simple Machine Learning Models
Emotion F1-score (Without Speech Features) F1-score (SpeechCueLLM)
Happy 59.73% 60.56% (+0.83%)
Sad 80.23% 82.21% (+1.98%)
Neutral 71.84% 73.75% (+1.91%)
Angry 65.54% 66.41% (+0.87%)
Excited 71.55% 73.71% (+2.16%)
Frustrated 68.25% 69.56% (+1.31%)
Table 9: F1-Scores for Different Emotions
[…]
angry+frustracted 86.02% 0.36% 9.25% 4.35%
happy_exc 0.75% 83.77% 13.51% 1.96%
neutral 14.86% 6.34% 74.80% 4.00%
sad 12.24% 0.41% 7.21% 80.14%
Table 10: Normalized Confusion Matrix - Without Speech Features
[…]
angry+frustracted 86.21% 0.73% 10.16% 2.90%
happy_exc 1.20% 82.52% 14.69% 1.58%
neutral 10.79% 5.74% 79.81% 3.57%
sad 10.76% 0.68% 8.17% 80.38%
Table 11: Normalized Confusion Matrix - SpeechCueLLM
To evaluate the effectiveness of our extracted speech features in capturing emotion-related information, we conducted experiments using various machine learning models on both numerical and categorical representations of these features. Tables 7 and 8 present the performance of Random Forest (RF), Support Vector Machine (SVM), and Multi-Layer Perceptron (MLP) models on the IEMOCAP and MELD datasets.